On another planet, the isotopes of titanium have t
Answer #1

Average atomic mass = 45.95263*0.746) + (47.94795*0.166)+ (49.94479 *0.088)= 46.635

---------------------------------------------------------------------------------------

-1 charge in an ion signifies that one electron is extra. Number of proton is 9.

----------------------------------------

Atomic number of Ba = 56. Ba^2+ has 56-2 = 54 electrons

--------------------------------------------------

there is 4 extra +ve ion (proton) in the ion . The charge is +4